# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v22

@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [256]. A feature model and a data set [^99] https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [121]Sandaruwan et al. [257] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [258] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [259] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [260] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [261] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [262] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [263] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [264] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [265]. The study by Shalinda and Munasinghe [266] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [267] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [268] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [269] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [122] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset100100100https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [270] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [271] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [272] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [273] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Wickramaarachchi et al. [274] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech.
@@
After the release of the SOLD dataset by Ranasinghe et al. [132] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [275] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [276] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [277] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Ranasinghe et al. [278] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task101101101https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [279] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [280] used mBERT to identify hate and offensive content in Sinhala.
@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [285, 286, 287, 288, 289] for Sinhala. For this, Arukgoda et al. [166] have proposed a system named Aruth based on the Lesk Algorithm[290]. An online tool102102102http://aruth.herokuapp.com/, an API103103103https://bit.ly/3sJEYbS of the algorithm, and code along with data on github104104104https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [291] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [292]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [293]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [294]. The study by Wimalasuriya [295], which has the same name as the above work by Welgama [294], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [296]. The study by Rathnayake et al. [297] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [298] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [141] introduced the XL-sum105105105https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website106106106https://www.bbc.com/sinhala. The M3LS107107107https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [142] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website108108108https://www.bbc.com/sinhala. Further, they claim that MT5 [299] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [301]Wickramasinghe et al. [302] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)109109109https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [135] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [303] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [327] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [328], active learning [329], and speaker adaptation [330]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [331] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [332]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [333]Priyadarshani [334] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [335], genetic algorithms [336], and syllable segmentation method utilizing acoustic envelopes [337]. The method proposed by Gunasekara and Meegama [338] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [339]. The work by Karunanayake et al. [340] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [341] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [342] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [343] and Subspace Gaussian Mixture Model (SGMM) [343] in Sinhala speech recognition. In the follow-up work, Gamage et al. [344] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [345] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [346]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result.  Karunathilaka et al. [347] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [348] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [323] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [349]. They have used the Kaldi speech recognition toolkit [350] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [351], proposed a Sinhala speech recognition system. The study by Azir et al. [352] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [353] is based on Tacotron [354] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [355], Nanayakkara and Weerasinghe [356] used DeepSpeech110110110https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [357] used an earlier version of the CMUSphinx toolkit111111111https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [358] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [359, 360] to automatically generate Sinhala subtitles from Speech.
@@ 3.17.3 Speech-to-Speech
Layansan et al. [361] created a speech-to-speech translation system for Sinhala on the Android platform. The system developed by Rajapakshe et al. [362] is also speech-to-speech in the sense that, it is a chatbot for scheduling medical appointments and giving medical advice where the front end contains speech recognition and voice synthesizer components that interfaced with a chatbot component in the back end.
@@ 3.18.1 Printed Text
The XTREME-UP112112112https://github.com/google-research/xtreme-up data set created by Ruder et al. [134] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [373] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [374, 375]. They then used hidden Markov model-based optimization on the recognized Sinhala script [376]. Similarly, Hewavitharana et al. [377] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [116], Herath and Medagoda [117] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [378], by Ajward et al. [379], and by Madushanka et al. [380].  Karunanayaka et al. [381] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [382] mainly involving the nearest-neighbor method [383, 384]. A study by Ediriweera [385] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [386] has been extended to be online and made available to use via desktops [387] and hand-held devices [388] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [389]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [390]Premachandra et al. [391] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [392]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [393].  Jayawickrama et al. [394] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [395] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [396]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [106] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [397].  Anuradha et al. [398] used Tesseract 3113113113https://tesseract-ocr.github.io/ [399] for Sinhala OCR. The later study by [400] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [401] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [402], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [317]. A study on Sinhala text extraction from social media images (memes) was conducted by Samarajeewa and Ranathunga [403]. They specifically handled the character-touching issue. The study by Walawage and Ranathunga [404] attempted to devise a feature set to separately identify Sinhala and English text on social media images (memes). The study by de Silva and Liyanage [405] used Convolutional Spiking Neural Networks to extract Sinhala text from YouTube thumbnails.  Chanda et al. [406] proposed a Gaussian kernel SVM based method for word-wise Sinhala, Tamil, and English script identification. The work by Vasantharajan et al. [107] adapted the Tesseract engine to handle non-Unicode (legacy fonts) in pdf documents to create a Tamil-Sinhala-English parallel corpus.

### v23

@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [266]. A feature model and a data set105105105https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [124]Sandaruwan et al. [267] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [268] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [269] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [270] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [271] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [272] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [273] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [274] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [275]. The study by Shalinda and Munasinghe [276] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [277] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [278] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [279] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [125] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset106106106https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [280] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [281] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [282] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [283] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Wickramaarachchi et al. [284] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech. The study by Muthuthanthri and Smith [285] used an annotated dataset from Facebook with TF-IDF and fastText embedding features with SVM, CNN, RNN, LSTM as well as BERT and GPT2. They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data.
@@
After the release of the SOLD dataset by Ranasinghe et al. [135] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [286] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [287] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [288] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Ranasinghe et al. [289] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task107107107https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [290] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [291] used mBERT to identify hate and offensive content in Sinhala.
@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [296, 297, 298, 299, 300] for Sinhala. For this, Arukgoda et al. [172] have proposed a system named Aruth based on the Lesk Algorithm[301]. An online tool108108108http://aruth.herokuapp.com/, an API109109109https://bit.ly/3sJEYbS of the algorithm, and code along with data on github110110110https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [302] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [303]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [304]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [305]. The study by Wimalasuriya [306], which has the same name as the above work by Welgama [305], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [307]. The study by Rathnayake et al. [308] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [309] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [144] introduced the XL-sum111111111https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website112112112https://www.bbc.com/sinhala. The M3LS113113113https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [145] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website114114114https://www.bbc.com/sinhala. Further, they claim that MT5 [310] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set. The first Sinhala multi-document summarizing dataset, M2DS115115115https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al. [146]. The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary. Thier code and pre-trained models are also available on github116116116https://github.com/KushanMH/m2ds.
@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [312]Wickramasinghe et al. [313] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)117117117https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [138] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [314] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [338] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [339], active learning [340], and speaker adaptation [341]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [342] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [343]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [344]Priyadarshani [345] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [346], genetic algorithms [347], and syllable segmentation method utilizing acoustic envelopes [348]. The method proposed by Gunasekara and Meegama [349] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [350]. The work by Karunanayake et al. [351] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [352] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [353] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [354] and Subspace Gaussian Mixture Model (SGMM) [354] in Sinhala speech recognition. In the follow-up work, Gamage et al. [355] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [356] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [357]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result. Later, they conducted further development on their model in a follow-up work [358].  Karunathilaka et al. [359] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [360] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [334] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [361]. They have used the Kaldi speech recognition toolkit [362] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [363], proposed a Sinhala speech recognition system. The study by Azir et al. [364] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [365] is based on Tacotron [366] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [367], Nanayakkara and Weerasinghe [368] used DeepSpeech118118118https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [369] used an earlier version of the CMUSphinx toolkit119119119https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [370] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [371, 372] to automatically generate Sinhala subtitles from Speech. The work by [373] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children.
@@ 3.17.3 Speech-to-Speech
Layansan et al. [374] created a speech-to-speech translation system for Sinhala on the Android platform. The system developed by Rajapakshe et al. [375] is also speech-to-speech in the sense that, it is a chatbot for scheduling medical appointments and giving medical advice where the front end contains speech recognition and voice synthesizer components that interfaced with a chatbot component in the back end. The work by Athas and Pirapuraj [376] uses Google APIs (speech-to-text API, text-to-text-translator API, and text-to-speech synthesizer API) to achieve voice translation for Sinhala and Tamil.
@@ 3.18.1 Printed Text
The XTREME-UP120120120https://github.com/google-research/xtreme-up data set created by Ruder et al. [137] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [387] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [388, 389]. They then used hidden Markov model-based optimization on the recognized Sinhala script [390]. Similarly, Hewavitharana et al. [391] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [119], Herath and Medagoda [120] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [392], by Ajward et al. [393], and by Madushanka et al. [394].  Karunanayaka et al. [395] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [396] mainly involving the nearest-neighbor method [397, 398]. A study by Ediriweera [399] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [400] has been extended to be online and made available to use via desktops [401] and hand-held devices [402] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [403]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [404]Premachandra et al. [405] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [406]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [407].  Jayawickrama et al. [408] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [409] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [410]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [109] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [411].  Anuradha et al. [412] used Tesseract 3121121121https://tesseract-ocr.github.io/ [413] for Sinhala OCR. The later study by [414] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [415] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [416], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. The work by Thamarasee and Surendra [417] used histogram-oriented gradient descriptor (HOG descriptor) [418] and SVMs to recognize 400+ variations of Sinhala characters.
An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [328]. A study on Sinhala text extraction from social media images (memes) was conducted by Samarajeewa and Ranathunga [419]. They specifically handled the character-touching issue. The study by Walawage and Ranathunga [420] attempted to devise a feature set to separately identify Sinhala and English text on social media images (memes). The study by de Silva and Liyanage [421] used Convolutional Spiking Neural Networks to extract Sinhala text from YouTube thumbnails.  Chanda et al. [422] proposed a Gaussian kernel SVM based method for word-wise Sinhala, Tamil, and English script identification. The work by Vasantharajan et al. [110] adapted the Tesseract engine to handle non-Unicode (legacy fonts) in pdf documents to create a Tamil-Sinhala-English parallel corpus.